OpenStudy (anonymous):

Given the function f(x) = 0.5(3)x, what is the value of f−1(7)? @shamil98

OpenStudy (anonymous):

@Salmon

OpenStudy (anonymous):

Let f(x)=y y=0.5(3)x=1.5x x=y/1.5 so f-1(x)=x/1.5 f-1(7)=7/1.5=14/3=4.66667

OpenStudy (anonymous):

0.565 1.140 1.771 2.402 these ARE the options @shamil98

OpenStudy (anonymous):

@Salmon

OpenStudy (anonymous):

i meant to tagg you

OpenStudy (anonymous):

@Salmon

OpenStudy (anonymous):

Its fine...even I am not getting a answer matching the options..

OpenStudy (anonymous):

ugh so what do I do?

OpenStudy (anonymous):

@Nurali @nincompoop Can you help..

OpenStudy (anonymous):

@Zarkon help!!!!

OpenStudy (rane):

@sania32